Four identical bricks are placed on top of each other at the edge of a table. Is it possible to slide them horizontally across each other in such a way that the projection of the topmost one is completely outside the table? What is the theoretical limit to the displacement of the topmost brick if the number of bricks is arbitrarily increased?


Grade:upto college level

1 Answers

ROSHAN MUJEEB
3 years ago

mx=m(l2−x)mx=m(l2-x)
x=l4x=l4
S=l4+l2=3l4